40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Sep 19 01:49:52 2020, Koji, Update, General, M4.5 EQ in LA 
    Reply  Sat Sep 19 11:27:04 2020, rana, Update, General, M4.5 EQ in LA Sep18-EQ.pdf
       Reply  Sat Sep 19 18:08:34 2020, rana, Update, General, M4.5 EQ in LA 
    Reply  Sat Sep 19 19:14:59 2020, Koji, Update, General, ITMX released / ETMY UR magnet knocked off IMG_6218.jpeg
       Reply  Sun Sep 20 11:41:54 2020, gautam, Update, General, IMC re-locked 
       Reply  Mon Sep 21 12:40:58 2020, gautam, Update, General, Workable IFO recovery IFOrecovery.pngIMG_5345.JPG
Message ID: 15590     Entry time: Mon Sep 21 12:40:58 2020     In reply to: 15585
Author: gautam 
Type: Update 
Category: General 
Subject: Workable IFO recovery 

See Attachment #1.

  • The original ETMY actuation matrix was the naive one so I simply retuned everything by adding the butterfly mode appropriately.
  • The cross coupling between the DOFs has not been characterized, but the local damping, Oplev loops, POX/POY locking, simpel Michelson locking, and ASS dither alignemnt all seem to work so I'm thinking this is good enough for now.
  • A copy of the ETMYaux.db file was made, and the slow bias voltage was redistributed among the three available face OSEMs - note that the magnet polarity has to be taken into account.
  • The series resistance on the slow path was reduced from 430 ohms, 1W to 100 ohms, 3W, to allow DC alignment of the cavity axis to the beam axis.
  • Vertex Oplevs were re-centered on their respective QPDs after locking POX/POY and running the ASS dither alignment.
  • The AS spot was a little low on the camera - presumably the SR2/SR3 got macroscopically misaligned. I re-centered the beam on the camera on the AS table (by moving the camera, the beam path was not disturbed).

The beam spot on ETMY looks weird (looks almost like a TEM10 mode), but the one on ITMY seems fine, see Attachment #2. Wonder what's going on there, maybe just a focusing problem?

Quote:
 

We need a vent to fix the suspension, but until then what we can do is to redistribute the POS/PIT/YAW actuations to the three coils.

Attachment 1: IFOrecovery.png  326 kB  Uploaded Mon Sep 21 13:41:16 2020  | Hide | Hide all
IFOrecovery.png
Attachment 2: IMG_5345.JPG  4.141 MB  Uploaded Mon Sep 21 13:47:58 2020  | Hide | Hide all
IMG_5345.JPG
ELOG V3.1.3-